# ADFormer: Aggregation Differential Transformer for Passenger Demand Forecasting

### v1

@@ 4 Experiments
Table 1: Performance comparison of different methods on public demnad datasets.
@@ 4.1.3 Model Parameters
The experimental environment is configured as follows: the Python version is 3.10.0, the CUDA version is 12.1, and the PyTorch version is 2.5.1. We conduct experiments on an NVIDIA GeForce RTX 3090 with 24GB of memory. The AdamW optimizer is used in model training, with an initial learning rate of 1e-3, decaying to 1e-4. We explore hidden dimension $\in{\{32,64,128\}}$, the depth of encoder $\in{\{4,6,8\}}$ and the impact of the number of spatial clusters and hierarchical levels in the Parameter Study.
@@ 4.5 Case Study
Table 3: The exploration of spatial aggregation numbers.
We visualize the high-level spatial correlations learned by our model to better illustrate the effectiveness of our approach. As illustrated in Figure 5, the urban area in the Xi’an dataset is divided into hexagonal grids. In (a), there are six clusters filled with dark or light red. In reality, these grids mostly contain residential areas. The second row of maps displays three grids that are relatively similar. These grids not only contain a large number of residential areas but also include major intersections, indicating their spatial similarity. Figure (b) shows larger clusters that are formed by aggregating the smaller clusters in (a), incorporating more regions. It merges the smaller potential clusters in (a) into broader clusters. Finally, Figure (c) not only presents the clustering results at the highest level but also illustrates the relationships between clusters, where grids with the same color belong to the same cluster. The core idea of spatial aggregation is to break the artificially defined spatial boundaries and view spatial correlations from a holistic perspective. The concept of temporal aggregation is similar to this.

### v2

@@ 4 Experiments
Table 1: Performance comparison of different methods on public demand datasets.
@@ 4.1.3 Model Parameters
The experimental environment is configured as follows: the Python version is 3.10.0, the CUDA version is 12.1, and the PyTorch version is 2.5.1. We conduct experiments on an NVIDIA GeForce RTX 3090 with 24GB of memory. The AdamW optimizer is used in model training, with an initial learning rate of 1e-3, decaying to 1e-4. We explore hidden dimension $\in{\{32,64,128\}}$, the depth of encoder $\in{\{4,6,8\}}$ and impact of number of spatial clusters and hierarchical levels in the Parameter Study.
@@ 4.5 Case Study
Table 3: Impact of spatial cluster numbers and hierarchical levels.
We visualize the high-level spatial correlations learned by our model to better illustrate the effectiveness of our approach. As illustrated in Figure 5, the urban area in the Xi’an dataset is divided into hexagonal grids. In (a), there are six clusters filled with dark or light red. In reality, these grids mostly contain residential areas. The second row of maps displays three grids that are relatively similar. These grids not only contain a large number of residential areas, but also include major intersections, indicating their spatial similarity and the potential for aggregation. Subfigure (b) shows larger clusters that are formed by aggregating the smaller clusters in (a), incorporating more regions. Finally, subfigure (c) presents the aggregation results at the highest level and illustrates the relationships between clusters, where grids of the same color belong to the same cluster. The core idea of spatial aggregation is to break artificially defined spatial boundaries and view spatial correlations from a holistic perspective. The concept of temporal aggregation is similar to this.